OpenStudy (anonymous):

Solving rational equations...need help with the concept, then I can do the individual questions myself.

OpenStudy (anonymous):

(7/8) - (1/x) = (3/4). Solve for x.

OpenStudy (anonymous):

I tried to get a common denominator of 8x, so I ended up with (7x/8x) - (8/8x) = (6x/8x). That became (7x - 8)/(8x) = (6x/8x), and then 56x^2 - 64x = 48x^2 became 8x^2 - 64x = 0.

OpenStudy (anonymous):

move the 7/8 over (−1/x)=(−7/8)+(6/8)=(−1/8)

OpenStudy (anonymous):

That's not how you solve this kind of problem. Algebra II. Fractions aren't in my answers, either. Thanks though.

OpenStudy (anonymous):

@linda3 Please help :)

OpenStudy (anonymous):

7/8- 1/x - 3/4 multiply by 8 7 - 8/x - 6 1 - 8/x make answer = 0 1 - 8/x = 0 -8/x = -1 8/x = 1 multiply by x so this leaves you with x = 8 -_-

OpenStudy (anonymous):

Okay...I just want a second opinion b/c that's not anything like how I was taught to solve a rational equation.

OpenStudy (anonymous):

@Kainui @Nurali @mathstudent55 @Australopithecus

OpenStudy (goformit100):

Sir/Ma'am, A Warm Welcome To Open Study. Sir/Ma'am please show your Working for the question you have posted.

OpenStudy (anonymous):

I already have.

OpenStudy (anonymous):

It's the second comment on this post. I just multiplied the first fraction by x/x, the second by 8/8, and the third by 2x/2x. Then I cross multiplied to get the aforementioned equation, and I converted it into standard form.

OpenStudy (anonymous):

@goformit100 Hello?

OpenStudy (anonymous):

@amistre64 @kmeis002 Please help.

OpenStudy (anonymous):

ok -_-

OpenStudy (anonymous):

What?

OpenStudy (kmeis002):

You can multiple by 8x for a common denominator. This should cancel out the x in second term and you should be left with something like: \[7x - 8 = 6x \] Then you may collect terms and solve from there.

OpenStudy (anonymous):

Okay $w3G, I guess you're right...Let's give it a try.

OpenStudy (anonymous):

┻━┻ ︵ヽ(`Д´)ノ︵ ┻━┻ lul

OpenStudy (amistre64):

(7/8) - (1/x) = (3/4) multiply thru by 8x to clear the fractions 7x - 8 = (3/4)8x 7x - 8 = 6x

OpenStudy (amistre64):

we could also notice that 3/4 = 6/8 and see that 7-1 = 6 if x=8 :)

OpenStudy (anonymous):

Great, thanks guys!!!

OpenStudy (anonymous):

@$w3G_Godd I have NO idea what that drawing is of.

OpenStudy (anonymous):

its me flipping 2 tables [̲̅$̲̅(̲̅5̲̅)̲̅$̲̅] ヾ(⌐■_■)ノ♪